# Self-Supervised Compression and Artifact Correction for Streaming Underwater Imaging Sonar

### v1

@@ Abstract
Real-time imaging sonar has become an important tool for underwater monitoring in environments where optical sensing is unreliable. Its broader use is constrained by two coupled challenges: highly limited uplink bandwidth and severe sonar-specific artifacts (speckle, motion blur, reverberation, acoustic shadows) that affect up to 98% of frames. We present SCOPE, a self-supervised framework that jointly performs compression and artifact correction without clean–noise pairs or synthetic assumptions. SCOPE combines (i) Adaptive Codebook Compression (ACC), which learns frequency-encoded latent representations tailored to sonar, with (ii) Frequency-Aware Multiscale Segmentation (FAMS), which decomposes frames into low-frequency structure and sparse high-frequency dynamics while suppressing rapidly fluctuating artifacts. A hedging training strategy further guides frequency-aware learning using low-pass proxy pairs generated without labels. Evaluated on months of in-situ ARIS sonar data, SCOPE achieves a structural similarity index (SSIM) of 0.77, representing a 40% improvement over prior self-supervised denoising baselines, at bitrates down to $\leq 0.0118$ bpp. It reduces uplink bandwidth by more than 80% while improving downstream detection. The system runs in real time, with 3.1 ms encoding on an embedded GPU and 97 ms full multi-layer decoding on the server end. SCOPE has been deployed for months in three Pacific Northwest rivers to support real-time salmon enumeration and environmental monitoring in the wild. Results demonstrate that learning frequency-structured latents enables practical, low-bitrate sonar streaming with preserved signal details under real-world deployment conditions.
@@ 1 Introduction
Imaging sonar has become increasingly important for real-time underwater monitoring in conditions where optical and infrared sensors are unreliable [jones2021adaptive, wei2022monitoring]. By transmitting acoustic pulses and reconstructing echoes into spatial imagery, sonar enables perception in turbid or low-light environments where cameras fail [hu2024underwater, shi2024sonar]. Recent advances in hardware affordability and wider access to 5G and satellite connectivity have expanded the feasibility of sonar-based analytics in remote settings, including offshore rescue [hu2024underwater], subsea infrastructure inspection [shi2024sonar], and seasonal fishery management [10.1145/3666025.3699323, kay2022caltechfishcountingdataset]. These applications highlight the broader societal value of imaging sonar in protecting human safety, supporting industrial operations, and enabling ecological monitoring [jahanbakht2021internet].
Despite its growing importance, streaming sonar reliably in uncontrolled, wild environments remains difficult. Limited infrastructure and unstable uplinks restrict transmission capacity [10228912], while the imagery itself is degraded by complex artifacts. For instance, in our field studies (see Suppl.), a single satellite uplink (e.g., Starlink) averaged 16.1 Mbps with a standard deviation of 5.7 and fluctuated between 1.1 and 39.9 Mbps. When shared across multiple observation points, the per-site bandwidth dropped to an average of 4.7 Mbps with a standard deviation of 1.9 and a range of 0 to 14 Mbps. This is well below the more than 24 Mbps required to stream compressed sonar video at full resolution (1146x2138@15 FPS), leaving a substantial portion of data untransmitted (see Suppl.). H.264/H.265 at the same CRF reach only $\sim$8.03 Mbps on standard video, showing sonar data is harder to compress (see Suppl.), due to pixel-wise noise fluctuations [Kwan]. Meanwhile, sonar images suffer from persistent speckle, motion blur, reverberations, and acoustic shadows [10274511, rs16152815]. Motion blur affects 98% of frames [kay2022caltechfishcountingdataset, jones2021adaptive], while reverberations and shadows can obscure objects by up to 6× their actual size [10.1145/3666025.3699323]. These conditions reduce perceptual quality, increase entropy, and hinder both compression efficiency and downstream analysis.
@@
To address these challenges, we present SCOPE, a real-time framework for Self-supervised COmpression and artifact COrrection in sonar Processing and strEaming. SCOPE is built on a Variational Autoencoder (VAE) backbone that performs latent-space compression, but extends it with mechanisms tailored for sonar. Standard VAEs can be inefficient because of large latent spaces, while generic codebook methods often collapse or fail to capture subtle patterns. We introduce Adaptive Codebook Compression (ACC) to learn frequency-encoded latent representations that stabilize training and preserve small, concealed objects. We further propose Frequency-Aware Multiscale Segmentation (FAMS) to separate low-frequency background structures from sparse high-frequency dynamics, suppressing artifacts in the process. Finally, a hedging training strategy leverages low-pass proxy pairs to guide frequency-aware learning without requiring clean–noise supervision. Together, these components enable SCOPE to jointly compress and correct sonar video, providing efficient low-bitrate transmission while retaining the details needed for reliable underwater monitoring.
@@
  • We develop and field-deploy SCOPE, a self-supervised framework that integrates compression and artifact correction to support real-time sonar streaming under practical constraints.
@@
Figure 1: SCOPE is a self-supervised framework for joint compression and artifact correction of image sonar streaming.
@@ 3 Background and Motivations
Sonar devices have become increasingly affordable, from under \$2,000 for handheld units [field2025] to \$25,000 for advanced systems like ARIS [joslin2019imaging]. Improved 5G and satellite connectivity enables broader real-time applications in remote areas, such as offshore rescue, disaster warnings, and fishery management. In turbid, low-light underwater environments where optical sensors fail, sonar provides a reliable alternative for rescues and fishery conservation [sonarrescue, 10.1145/3666025.3699323], highlighting its social impact in emergency response and environmental monitoring.
Challenges from wild environments. Despite wider availability, deploying sonar in remote areas faces hurdles from limited infrastructure. Terrestrial networks cover only 15% of the Earth [cellularcoverage], while satellite links suffer from high latency (up to 600 ms) and unstable uplink bandwidth, especially in dense vegetation, mountains, or harsh weather [10.1145/3664647.3680785]. These constraints hinder real-time sonar transmission. For example, an ARIS Explorer 3000 [aris3000] at 1.8MHz outputs 1280×800@15 FPS, a downsample from higher native resolution used in previous deployments that degrades performance, still requiring $\geq$10.7 Mbps after codecs, whereas typical satellite uplinks provide only 4.7 Mbps (47% of the required bandwidth), making compression essential.
Artifacts of sonar data. Sonar data from wild environments often contains artifacts such as speckle noise (high-frequency granular patterns), motion blur, acoustic shadows, and multipath reverberations (Fig. 2) that obscure object boundaries, reduce contrast, lower perceptual quality, increase entropy, hinder compression, and lower analytic accuracy in tasks such as detection, tracking, and segmentation. Effective correction is essential for maintaining signal fidelity and reliable sonar imagery.
@@ 4 Method
We develop SCOPE, a self-supervised real-time sonar streaming framework that jointly compresses data and corrects artifacts without requiring clean-noise pairs or synthetic data, and is deployed across three rivers (Fig. 1).
@@ 4.1 SCOPE Overview
As shown in Fig. 4, a noisy input $I_{i}$ is encoded into a latent representation $\hat{I}_{i}$, which is processed by Adaptive Codebook Compression (ACC) and Frequency-Aware Multiscale Segmentation (FAMS). ACC employs a VQ-VAE-based lossy LIC scheme to enhance representation learning of characteristic frequency components tailored for underwater sonar data, producing $k$ index maps for hierarchical sonar compression. FAMS performs soft frequency segmentation to isolate high-frequency components $H_{i}$ (e.g., fish), generating $k$ layers per frame by decoding the index maps $k$ times (Fig. 4, upper blue region). Compression is achieved by transmitting only $H_{i}$ for each frame, while the static low-frequency background $S$ is transmitted weekly. At the receiver, $S$ and $H_{i}$ are recombined to reconstruct a clean, artifact-free frame. Training is guided by a self-supervised hedging strategy (Sec. 4.4).
Figure 5: Pixel-wise temporal analysis of intensity fluctuations across sonar frames.
@@ 4.3 Frequency-Aware Multiscale Segmentation
Fig.5 shows a pixel-wise change heatmap across frames in real-world underwater sonar video. Brighter areas (7.73% of total), from white to yellow, indicate high-frequency object motions like fish and prawns (red box in Fig.3), while darker areas, from red to black, correspond to low-frequency static structures (blue box). The 2nd and 4th waveforms represent artifacts such as speckle noise, motion blur, and acoustic shadows. Comparing these waveforms reveals clear distinctions between high-frequency signals $H_{i}$, low-frequency static signals $S$, and artifacts.
@@ 4.4 Hedging Training Strategy
Low-pass proxy pair. As shown in Fig. 8(b)(c), we generate low-pass proxy pairs by first applying the Mixture of Gaussians (MOG) method for change detection to highlight dynamic elements (e.g., fish), followed by a guided filter acting as a low-pass filter. This process coarsely segments temporal frequency components, preserving low-frequency structures and edges while blurring static backgrounds, thereby emphasizing dynamic regions for segmentation guidance. This approach suits underwater sonar data, where dynamic objects show higher temporal variance captured efficiently by MOG, and the guided filter preserves low-frequency spatial structures for robust feature learning. Combining temporal and spatial filtering provides a clear supervisory signal that emphasizes salient motion patterns while suppressing noise and static clutter, enabling effective learning of meaningful frequency representations.
@@ 5.1 Experimental Settings
Figure 9: Visual comparison of self-supervised denoising methods on Underwater Sonar Imaging Dataset.
@@ 5.2 Evaluation of Artifact Correction
We target diverse artifact types rather than treating them as general noise, positioning denoising as a coarse correction step tailored to sonar data. Our framework, SCOPE, is compared with state-of-the-art self-supervised denoising methods for real-world scenarios, including AP-BSN [9878719], LG-BPN [10204077], SDAP [10377346], and DDPG [Garber_2024_CVPR].
@@
Table 1: Comparison of denoising methods on a real-world underwater sonar dataset.
@@ 5.3 Evaluation of Compression
Table 2: Comparisons of different compression models.
@@ 5.5 Ablation Study
Frequency-Aware Multiscale Segmentation (FAMS). The effect of proposed FAMS is illustrated in Figs. 4 and 7. It suppresses high-frequency artifacts while segmenting low-frequency background structures and high-frequency object signals (e.g., fish). This method extends beyond sonar, showing potential applicability in imaging domains such as MRI and remote sensing. In MRI, it could contribute to improved separation of tissue types. In remote sensing, it could contribute to more accurate change detection and geospatial monitoring.
@@ 6 Conclusion
In this paper, we presented a self-supervised SCOPE framework for underwater imaging sonar compression and artifact correction. By integrating ACC and FAMS with hedging training, it suppressed artifacts while preserving critical high-frequency signals. Experiments showed superior compression ($\leq 0.0118$ bpp) suitable for 4.7 Mbps uplinks and artifact correction (SSIM: 0.7711), which improved downstream task accuracy. With 3.1 ms encoding and 97 ms total decoding for all layers, SCOPE enables real-time processing and has been deployed at three Pacific Northwest river sites for salmon and environmental monitoring, enabling practical sonar streaming. The method also shows promise for other imaging fields like MRI and remote sensing.

### v2

@@ Abstract
Real-time imaging sonar is crucial for underwater monitoring where optical sensing fails, but its use is limited by low uplink bandwidth and severe sonar-specific artifacts (speckle, motion blur, reverberation, acoustic shadows) affecting up to 98% of frames. We present SCOPE, a self-supervised framework that jointly performs compression and artifact correction without clean–noise pairs or synthetic assumptions. SCOPE combines (i) Adaptive Codebook Compression (ACC), which learns frequency-encoded latent representations tailored to imaging sonar, with (ii) Frequency-Aware Multiscale Segmentation (FAMS), which decomposes frames into low-frequency structure and sparse high-frequency dynamics while suppressing rapidly fluctuating artifacts. A hedging training strategy further guides frequency-aware learning using low-pass proxy pairs generated without labels. Evaluated on months of in-situ ARIS sonar data, SCOPE achieves a structural similarity index (SSIM) of 0.77, representing a 40% improvement over prior self-supervised denoising baselines, at bitrates down to $\leq 0.0118$ bpp. It reduces uplink bandwidth by more than 80% while improving downstream detection. The system runs in real time, with 3.1 ms encoding on an embedded GPU and 97 ms full multi-layer decoding on the server end. SCOPE has been deployed for months in three Pacific Northwest rivers to support real-time salmon enumeration and environmental monitoring in the wild. Results demonstrate that learning frequency-structured latents enables practical, low-bitrate sonar streaming with preserved signal details under real-world deployment conditions.
@@ 1 Introduction
Traditional sonar operated at high power for longer range and wider coverage, but had limited resolution, making it difficult to identify small targets and resolve fine details of objects. In contrast, modern imaging sonars provide high-resolution imagery at lower power, enabling detailed observation of underwater scenes and real-time monitoring in conditions where optical and infrared sensors are unreliable [jones2021adaptive, wei2022monitoring]. By transmitting acoustic pulses and reconstructing echoes into spatial imagery, sonar allows perception in turbid or low-light environments where cameras fail [hu2024underwater, shi2024sonar]. Advances in hardware affordability and access to 5G and satellite connectivity have expanded sonar-based analytics in remote settings, including offshore rescue [hu2024underwater], subsea infrastructure inspection [shi2024sonar], and seasonal fishery management [10.1145/3666025.3699323, kay2022caltechfishcountingdataset], highlighting its societal value in safety, industrial support, and ecological monitoring [jahanbakht2021internet].
Despite its growing importance, it remains difficult to reliably perform streaming imaging sonar in wild, uncontrolled environments. Limited infrastructure and unstable uplinks restrict transmission capacity [10228912], while the imagery itself is degraded by complex artifacts. For instance, in our field studies (see Suppl.), a single satellite uplink (e.g., Starlink) averaged 16.1 Mbps with a standard deviation of 5.7 and fluctuated between 1.1 and 39.9 Mbps. When shared across multiple observation points, the per-site bandwidth dropped to an average of 4.7 Mbps with a standard deviation of 1.9 and a range of 0 to 14 Mbps. This is well below the more than 24 Mbps required to stream compressed sonar video at full resolution (1146x2138@15 FPS), leaving a substantial portion of data untransmitted (see Suppl.). H.264/H.265 at the same CRF reach only $\sim$8.03 Mbps on standard video, showing imaging sonar data is harder to compress (see Suppl.), due to pixel-wise noise fluctuations [Kwan]. Meanwhile, imaging sonar suffers from persistent speckle, motion blur, reverberations, and acoustic shadows [10274511, rs16152815]. Motion blur affects 98% of frames [kay2022caltechfishcountingdataset, jones2021adaptive], while reverberations and shadows can obscure objects by up to 6× their actual size [10.1145/3666025.3699323]. These conditions reduce perceptual quality, increase entropy, and hinder both compression efficiency and downstream analysis.
@@
To address these challenges, we present SCOPE, a real-time framework for Self-supervised COmpression and artifact COrrection in imaging sonar Processing and strEaming. SCOPE is built on a Variational Autoencoder (VAE) backbone that performs latent-space compression, but extends it with mechanisms tailored for imaging sonar. Standard VAEs can be inefficient because of large latent spaces, while generic codebook methods often collapse or fail to capture subtle patterns. We introduce Adaptive Codebook Compression (ACC) to learn frequency-encoded latent representations that stabilize training and preserve small, concealed objects. We further propose Frequency-Aware Multiscale Segmentation (FAMS) to separate low-frequency background structures from sparse high-frequency dynamics, suppressing artifacts in the process. Finally, a hedging training strategy leverages low-pass proxy pairs to guide frequency-aware learning without requiring clean–noise supervision. Together, these components enable SCOPE to jointly compress and correct sonar video, providing efficient low-bitrate transmission while retaining the details needed for reliable underwater monitoring.
@@
  • We develop and field-deploy SCOPE, a self-supervised framework that integrates compression and artifact correction to support real-time imaging sonar streaming under practical constraints.
@@
Figure 1: SCOPE is a self-supervised framework for joint compression and artifact correction of imaging sonar streaming.
@@ 3 Background and Motivations
Sonar devices have become increasingly affordable, from under \$2,000 for handheld units [field2025] to \$25,000 for advanced imaging sonar systems like ARIS [joslin2019imaging], which are cheaper and higher-resolution than traditional sonar. Improved 5G and satellite connectivity enables broader real-time applications in remote areas, such as offshore rescue, disaster warnings, and fishery management. In turbid, low-light underwater environments where optical sensors fail, sonar provides a reliable alternative for rescues and fishery conservation [sonarrescue, 10.1145/3666025.3699323], highlighting its social impact in emergency response and environmental monitoring.
Challenges from wild environments. Despite wider availability, deploying sonar in remote areas faces hurdles from limited infrastructure. Terrestrial networks cover only 15% of the Earth [cellularcoverage], while satellite links suffer from high latency (up to 600 ms) and unstable uplink bandwidth, especially in dense vegetation, mountains, or harsh weather [10.1145/3664647.3680785]. These constraints hinder real-time imaging sonar transmission. E.g., an ARIS Explorer 3000 [aris3000] at 1.8MHz outputs 1280×800@15 FPS, a downsample from higher native resolution used in previous deployments that degrades performance, still requiring $\geq$10.7 Mbps after codecs, whereas typical satellite uplinks provide only 4.7 Mbps (47% of the required bandwidth), making compression essential.
Artifacts of sonar data. Imaging sonar data from wild environments often contains artifacts such as speckle noise (high-frequency granular patterns), motion blur, acoustic shadows, and multipath reverberations (Fig. 2) that obscure object boundaries, reduce contrast, lower perceptual quality, increase entropy, hinder compression, and lower analytic accuracy in tasks such as detection, tracking, and segmentation. Effective correction is essential for maintaining signal fidelity and reliable sonar imagery.
@@ 4 Method
We develop SCOPE, a self-supervised real-time imaging sonar streaming framework that jointly compresses data and corrects artifacts without requiring clean-noise pairs or synthetic data, and is deployed across three rivers (Fig. 1).
@@ 4.1 SCOPE Overview
As shown in Fig. 4, a noisy input $I_{i}$ is encoded into a latent representation $\hat{I}_{i}$, which is processed by Adaptive Codebook Compression (ACC) and Frequency-Aware Multiscale Segmentation (FAMS). ACC employs a VQ-VAE-based lossy LIC scheme to enhance representation learning of characteristic frequency components tailored for underwater imaging sonar data, producing $k$ index maps for hierarchical sonar compression. FAMS performs soft frequency segmentation to isolate high-frequency components $H_{i}$ (e.g., fish), generating $k$ layers per frame by decoding the index maps $k$ times (Fig. 4, upper blue region). Compression is achieved by transmitting only $H_{i}$ for each frame, while the static low-frequency background $S$ is transmitted weekly. At the receiver, $S$ and $H_{i}$ are recombined to reconstruct a clean, artifact-free frame. Training is guided by a self-supervised hedging strategy (Sec. 4.4).
Figure 5: Pixel-wise temporal analysis of intensity fluctuations across imaging sonar frames.
@@ 4.3 Frequency-Aware Multiscale Segmentation
Fig.5 shows a pixel-wise change heatmap across frames in real-world underwater imaging sonar video. Brighter areas (7.73% of total), from white to yellow, indicate high-frequency object motions like fish and prawns (red box in Fig.3), while darker areas, from red to black, correspond to low-frequency static structures (blue box). The 2nd and 4th waveforms represent artifacts such as speckle noise, motion blur, and acoustic shadows. Comparing these waveforms reveals clear distinctions between high-frequency signals $H_{i}$, low-frequency static signals $S$, and artifacts.
@@ 4.4 Hedging Training Strategy
Low-pass proxy pair. As shown in Fig. 8(b)(c), we generate low-pass proxy pairs by first applying the Mixture of Gaussians (MOG) method for change detection to highlight dynamic elements (e.g., fish), followed by a guided filter acting as a low-pass filter. This process coarsely segments temporal frequency components, preserving low-frequency structures and edges while blurring static backgrounds, thereby emphasizing dynamic regions for segmentation guidance. This approach suits underwater imaging sonar data, where dynamic objects show higher temporal variance captured efficiently by MOG, and the guided filter preserves low-frequency spatial structures for robust feature learning. Combining temporal and spatial filtering provides a clear supervisory signal that emphasizes salient motion patterns while suppressing noise and static clutter, enabling effective learning of meaningful frequency representations.
@@ 5.1 Experimental Settings
Figure 9: Visual comparison of self-supervised denoising methods on Underwater Imaging Sonar Dataset.
@@ 5.2 Evaluation of Artifact Correction
We target diverse artifact types rather than treating them as general noise, positioning denoising as a coarse correction step tailored to imaging sonar data. Our framework, SCOPE, is compared with state-of-the-art self-supervised denoising methods for real-world scenarios, including AP-BSN [9878719], LG-BPN [10204077], SDAP [10377346], and DDPG [Garber_2024_CVPR].
Table 1: Denoising methods comparison on real-world sonar data.
@@ 5.3 Evaluation of Compression
Table 2: Compression methods comparison on imaging sonar data.
@@ 5.5 Ablation Study
Frequency-Aware Multiscale Segmentation (FAMS). The effect of proposed FAMS is illustrated in Figs. 4 and 7. It suppresses high-frequency artifacts while segmenting low-frequency background structures and high-frequency object signals (e.g., fish). This method extends beyond imaging sonar, showing potential applicability in imaging domains such as MRI and remote sensing. In MRI, it could contribute to improved separation of tissue types. In remote sensing, it could contribute to more accurate change detection and geospatial monitoring.
@@ 6 Conclusion
In this paper, we presented a self-supervised SCOPE framework for underwater imaging sonar compression and artifact correction. By integrating ACC and FAMS with hedging training, it suppressed artifacts while preserving critical high-frequency signals. Experiments showed superior compression ($\leq 0.0118$ bpp) suitable for 4.7 Mbps uplinks and artifact correction (SSIM: 0.7711), which improved downstream task accuracy. With 3.1 ms encoding and 97 ms total decoding for all layers, SCOPE enables real-time processing and has been deployed at three Pacific Northwest river sites for salmon and environmental monitoring, enabling practical sonar streaming. The method also shows promise for other imaging fields such as MRI and remote sensing.
Acknowledgment
This research is supported by an NSERC Discovery Grant, a British Columbia Salmon Recovery and Innovation Fund (BCSRIF_2022_401), and a Mitacs Accelerate Cluster Grant. It also received additional funding from Experiment.com. Chi Xu’s work is supported by an NSERC Canada Graduate Scholarship–Doctoral (CGS-D) and a Mitacs Globalink Research Award. We are grateful for the trust and collaboration of the Pacific Salmon Foundation, Wild Salmon Center, the Heiltsuk, Haida, Kitasoo Xai’xais, Taku River Tlingit, and Gitga’at First Nations, as well as the Skeena Fisheries Commission and Gitanyow Fisheries Authority for their ongoing partnership in this work.